The chance of getting a particular disease sometime during your lifetime is estimated at 1 in 20.
Smoking increases the risk of developing the disease during your lifetime by 10%.
How many people out of a group of 1000 smokers would you expect to get the disease sometime during their lifetime?

Answer by stanbon(75887)

The chance of getting a particular disease sometime during your lifetime is estimated at 1 in 20.
Smoking increases the risk of developing the disease during your lifetime by 10%.
----
(1/20)+(0.10) = 0.05+0.10 = 0.15
-----------------------------------------
How many people out of a group of 1000 smokers would you expect to get the disease sometime during their lifetime?
---
Ans: 0.15*1000 = 150
=========================
Cheers,
Stan H.